OpenStudy (anonymous):

the function f(x)=2x^3-3x+5. Find the average rate of change between points were x=1 and x=2

OpenStudy (anonymous):

\[f[x]=2 x^3-3 x+5 \]\[f \text{ ' }[x]=6 x^2-3 \]\[\frac{f\text{ ' }[2]-f\text{ ' }[1]}{2-1}=18 \]Hope I am right on the formulation of this one.

OpenStudy (anonymous):

Plot attached.